# A Survey on Large Language Models for Code Generation

### v1

@@ 4.3.1. Model Architectures
Since the inception of the Transformer architecture for machine translation (Vaswani et al., 2017), it has become the de facto backbone for a multitude of large language models (LLMs) that address a wide range of downstream tasks. The Transformer and its derivatives owe their prominence to their exceptional ability to parallelize computation and their powerful representational capacities (Zhao et al., 2023b; Yoo et al., 2024). Through innovative scaling techniques, such as Mixture-of-Experts (MoE) (Shazeer et al., 2017; Cai et al., 2024) and Depth-Up-Scaling (DUS) (Kim et al., 2023), the capacity of Transformer-based LLMs has expanded to encompass hundreds of billions or even trillions of parameters. These scaled-up models have exhibited a range of emergent abilities (Kaplan et al., 2020; Hoffmann et al., 2022; Wei et al., 2022a), such as instruction following (Ouyang et al., 2022), in-context learning (Dong et al., 2022), and step-by-step reasoning (Wei et al., 2022b; Huang and Chang, 2022) that were previously unforeseen.
@@
Table 4. The overview of large language models (LLMs) with encoder-decoder architectures for code generation.
Architecture Model Institution Size Vocabulary Context Window Date Open Source
Encoder-Decoder PyMT5(Clement et al., 2020) Microsoft 374M 50K 1024+1024 2020-10
@@
Table 5. The overview of large language models (LLMs) with decoder-only architectures for code generation.
Architecture Model Institution Size Vocabulary Context Window Date Open Source
Decoder-Only GPT-C (Svyatkovskiy et al., 2020) Microsoft 366M 60K 1024 2020-05
@@ 4.3.2. Pre-training Tasks
where $\mathbf{x}_{\leq k}$ is the source sequence input and $\mathbf{x}_{<i}$ denotes the target sequence autoregressively generated so far. During the inference phase, pre-trained LLMs that have been trained on large-scale code corpus can generate code in a zero-shot manner without the need for fine-tuning. This is achieved through the technique of prompt engineering, which guides the model to produce the desired output [^11] For more information on prompt engineering, visit https://www.promptingguide.ai (Radford et al., 2019; Brown et al., 2020). Additionally, recent studies have explored the use of few-shot learning, also referred to as in-context learning, to enhance model performance further (Li et al., 2023d; Patel et al., 2023).
@@ 4.4. Instruction Tuning
After pre-training Large Language Models (LLM) on large-scale datasets, the next phase typically involves augmenting the model’s ability to process and follow instructions, known as instruction tuning. Instruction tuning generally refers to the supervised fine-tuning of pre-trained LLMs using datasets comprised of structured examples framed as natural language instructions (Wei et al., 2021; Ouyang et al., 2022; Iyer et al., 2022; Zhang et al., 2023d). Two exemplars of instruction data sampled from Code Alpaca (Chaudhary, 2023) are demonstrated in Figure 5. It capitalizes on the heterogeneity of instruction types, positioning instruction tuning as a form of multi-task prompted training that significantly enhances the model’s generalization to unseen tasks (Wei et al., 2021; Sanh et al., 2022; Ouyang et al., 2022; Chung et al., 2024).
In the realm of code generation, natural language descriptions serve as the instructions guiding the model to generate corresponding code snippets. Consequently, a line of research on instruction tuning LLMs for code generation has garnered substantial interest across academia and industry. To perform instruction tuning, instruction data are typically compiled from source code with permissive licenses (Husain et al., 2019; Kocetkov et al., 2022; Lozhkov et al., 2024) (refer to Section 4.1.2) or are constructed from synthetic code data (Luo et al., 2023; Wei et al., 2023; Yuxiang Wei, 2024) (refer to Section 4.2). These datasets are then utilized to fine-tune LLMs through a supervised learning paradigm. However, the substantial computational resources required for full parameter fine-tuning (FFT) LLM pose a notable challenge, particularly in scenarios with constrained resources (Ding et al., 2022a; Lialin et al., 2023). To mitigate this issue, parameter-efficient fine-tuning (PEFT) has emerged as a compelling alternative strategy, gaining increasing attention for its potential to reduce resource consumption (Ding et al., 2022a). In the following subsection, we categorize existing works based on their instruction-tuning strategies to provide a comprehensive and systematic review.
@@ 4.4.1. Full Parameter Fine-tuning
Full parameter fine-tuning (FFT) involves updating all parameters within a pre-trained model, as shown in Figure 6(a). This approach is often preferred when ample computational resources and substantial training data are available, as it typically leads to better performance. (Wang et al., 2021) introduces an encoder-decoder pre-trained language model for code generation, named CodeT5+. They instruction-tune this model on a dataset comprising 20k instruction samples from Code Alpaca (Chaudhary, 2023), resulting in an instruction-following model called InstructCodeT5+, which exhibited improved capabilities in code generation. (Luo et al., 2023) leverages the Evol-Instruct data synthesis technique from WizardLM (Xu et al., 2023) to evolve 20K code Alpaca (Chaudhary, 2023) instruction samples into a 78K code instruction dataset. This enriched dataset is then used to fine-tune the StarCoder base model, resulting in WizardCoder, which showcases notable advancements in code generation. In a similar vein, inspired by the successes of WizardCoder (Luo et al., 2023) and RRHF (Yuan et al., 2023), Pangu-Coder 2 (Shen et al., 2023) applies the Evol-Instruct method to generate 68k high-quality instruction samples from the initial 20k Code Alpaca (Chaudhary, 2023) instruction samples. Additionally, they introduces a novel reinforcement learning via Rank Responses to align Test & Teacher Feedback (RRTF), which further enhances the performance of Pangu-Coder 2 in code generation. Diverging from synthetic instruction data generation methods, OctoPack (Muennighoff et al., 2023) utilizes real-world data by curating CommitPack from the natural structure of Git commits, which inherently pair code changes with human-written instructions. This dataset, consisting of 4 terabytes of Git commits across 350 programming languages, is employed to fine-tune StarCoder (Li et al., 2023a) and CodeGeeX2 (Zheng et al., 2023c), leading to the instruction-following code models of OctoCoder and OctoGeeX for code generation, respectively. The most recent innovation comes from Magicoder (Wei et al., 2023), who proposes OSS-INSTRUCT, a novel data synthesis method that leverages open-source code snippets to generate high-quality instruction data for code generation. This approach seeks to reduce the bias often present in synthetic data generated by LLM. In line with OSS-INSTRUCT, the BigCode team introduces StarCoder2-15B-instruct (Yuxiang Wei, 2024), which they claim to be the first entirely self-aligned Large Language Model (LLM) for code generation, trained with a fully permissive and transparent pipeline. Moreover, (CodeGemma Team et al., 2024) harnesses open-source mathematics datasets, such as MATH (Hendrycks et al., 2021) and GSM8k (Cobbe et al., 2021), along with synthetically generated code following the OSS-INSTRUCT (Wei et al., 2023) paradigm, to instruction-tune CodeGemma 7B, yielding exceptional results in mathematical reasoning and code generation tasks.
@@ 4.4.2. Parameter-Efficient Fine-tuning
To mitigate the extensive computational and resource demands inherent in fine-tuning large language models (LLMs), the concept of parameter-efficient fine-tuning (PEFT) has emerged to focus on updating a minimal subset of parameters, which may either be a selection of the model’s parameters or an array of additional parameters specifically introduced for the tuning process (Ding et al., 2022a; Lialin et al., 2023). The categorization of these methods is depicted in Figure 6(b), (c), and (d). A plethora of innovative PEFT approaches have been developed, among which BitFit (Zaken et al., 2021), Adapter (Houlsby et al., 2019), Prompt tuning (Lester et al., 2021), Prefix-tuning (Li and Liang, 2021), LoRA (Hu et al., 2021), IA3 (Liu et al., 2022), QLoRA (Dettmers et al., 2024), and AdaLoRA (Zhang et al., 2023a) are particularly noteworthy. A seminal study in this field, LoRA (Hu et al., 2021), proposes a parameter update mechanism for a pre-trained weight matrix — such as those found in the key or value projection matrices of a Transformer block’s multi-head self-attention layer — by factorizing the update into two low-rank matrices. Crucially, all original model parameters remain frozen, with only the pair of low-rank matrices being trainable. After fine-tuning, the product of these low-rank matrices can be seamlessly incorporated into the existing weight matrix through an element-wise addition. This process can be formally described as:
@@
Despite the advancements in PEFT methods, their application in code generation remains limited. For instance, (Jiang and Kim, 2023) pioneered the use of parameter-efficient instruction-tuning on a Llama 2 (Touvron et al., 2023b) model with a single RTX 3090 GPU, leading to the development of a multilingual code generation model called CodeUp. More recently, ASTRAIOS (Zhuo et al., 2024) conducted a thorough empirical examination of parameter-efficient instruction tuning for code comprehension and generation tasks. This study yielded several perceptive observations and conclusions, contributing valuable insights to the domain.
@@ 4.5. Reinforcement Learning with Feedback
Large language models (LLMs) have exhibited remarkable instruction-following capabilities through instruction tuning. However, they often produce outputs that are unexpected, toxic, biased, or hallucinated outputs that do not align with users’ intentions or preferences (Ouyang et al., 2022; Wang et al., 2023c; Ji et al., 2023). Consequently, aligning LLMs with human preference has emerged as a pivotal area of research. A notable work is InstructGPT (Ouyang et al., 2022), which further fine-tunes an instruction-tuned model utilizing reinforcement learning with human feedback (RLHF) on a dataset where labelers have ranked model outputs in order of quality, from best to worst. This method has been instrumental in the development of advanced conversational language models, such as ChatGPT (OpenAI, 2022) and Bard (Manyika and Hsiao, 2023). Despite its success, acquiring high-quality human preference ranking data is a resource-intensive process (Lee et al., 2023). To address this, Reinforcement Learning from AI Feedback (RLAIF) (Bai et al., 2022; Lee et al., 2023) has been proposed to leverage powerful off-the-shelf LLMs (e.g., ChatGPT (OpenAI, 2022) and GPT-4 (Achiam et al., 2023)) to simulate human annotators by generating preference data.
@@
Building on RLHF’s success, researchers have explored reinforcement learning with feedback to enhance code generation in LLMs. Unlike RLHF, which relies on human feedback, this approach employs compilers or interpreters to automatically provide feedback on code samples through code execution on unit test cases, catalyzing the advancement of this research domain. CodeRL (Le et al., 2022) introduced an actor-critic reinforcement learning framework for code generation. In this setup, the language model serves as the actor-network, while a token-level functional correctness reward predictor acts as the critic. Generated code is assessed through unit test signals from a compiler, which can indicate compiler errors, runtime errors, unit test failures, or passes. CompCoder (Wang et al., 2022a) enhances code compilability by employing compiler feedback, including language model fine-tuning, compilability reinforcement, and compilability discrimination strategies. Subsequently, PPOCoder (Shojaee et al., 2023) integrates pre-trained code model CodeT5 (Wang et al., 2021) with Proximal Policy Optimization (PPO) (Schulman et al., 2017). This integration not only utilizes execution (i.e., compilers or interpreters) feedback to assess syntactic and functional correctness but also incorporates a reward function that evaluates the syntactic and semantic congruence between abstract syntax tree (AST) sub-trees and data flow graph (DFG) edges in the generated code against the ground truth. Additionally, the framework applies a KL-divergence penalty to maintain fidelity between the actively learned policy and the referenced pre-trained model, enhancing the optimization process. More recently, RLTF (Liu et al., 2023d) has proposed an online reinforcement learning framework that provides fine-grained feedback based on compiler error information and location, along with adaptive feedback that considers the ratio of passed test cases.
@@
Despite these successes, reinforcement learning algorithms face inherent limitations such as inefficiency, instability, extensive resource requirements, and complex hyperparameter tuning, which can impede the performance and scalability of LLMs. To overcome these challenges, recent studies have introduced various variants of RL methods that do not rely on PPO, including DPO (Rafailov et al., 2024), RRHF (Yuan et al., 2023), and sDPO (Kim et al., 2024). In essence, these methods aim to maximize the likelihood between the logarithm of conditional probabilities of preferred and rejected responses, which may be produced by LLMs with varying capabilities. Inspired by RRHF (Yuan et al., 2023), PanGu-Coder 2 (Shen et al., 2023) leverages a novel framework, Reinforcement Learning via Rank Responses to align Test & Teacher Feedback (RRTF), significantly enhancing code generation capabilities, as evidenced by pass@1 of 62.20% on the HumanEval benchmark.

### v2

@@ 5.3.1. Model Architectures
Since the inception of the Transformer architecture for machine translation (Vaswani et al., 2017), it has become the de facto backbone for a multitude of LLMs that address a wide range of downstream tasks. The Transformer and its derivatives owe their prominence to their exceptional ability to parallelize computation and their powerful representational capacities (Zhao et al., 2023b; Yoo et al., 2024). Through innovative scaling techniques, such as Mixture-of-Experts (MoE) (Shazeer et al., 2017; Cai et al., 2024a) and Depth-Up-Scaling (DUS) (Kim et al., 2023), the capacity of Transformer-based LLMs has expanded to encompass hundreds of billions or even trillions of parameters. These scaled-up models have exhibited a range of emergent abilities (Kaplan et al., 2020; Hoffmann et al., 2022; Wei et al., 2022a), such as instruction following (Ouyang et al., 2022), in-context learning (Dong et al., 2022), and step-by-step reasoning (Wei et al., 2022b; Huang and Chang, 2022) that were previously unforeseen.
@@
Table 7. The overview of LLMs with encoder-decoder architectures for code generation.
Model Institution Size Vocabulary Context Window Date Open Source
PyMT5(Clement et al., 2020) Microsoft 374M 50K 1024+1024 2020-10
@@
Table 8. The overview of LLMs with decoder-only architectures for code generation.
Model Institution Size Vocabulary Context Window Date Open Source
GPT-C (Svyatkovskiy et al., 2020) Microsoft 366M 60K 1024 2020-05
@@ 5.3.2. Pre-training Tasks
where $\mathbf{x}_{\leq k}$ is the source sequence input and $\mathbf{x}_{<i}$ denotes the target sequence autoregressively generated so far. During the inference phase, pre-trained LLMs that have been trained on large-scale code corpus can generate code in a zero-shot manner without the need for fine-tuning. This is achieved through the technique of prompt engineering, which guides the model to produce the desired output [^11] For more information on prompt engineering, visit https://www.promptingguide.ai (Radford et al., 2019; Brown, 2020). Additionally, recent studies have explored the use of few-shot learning, also referred to as in-context learning, to enhance model performance further (Li et al., 2023d; Patel et al., 2023).
@@
Figure 9. Comparison of instruction tuning with various fine-tuning strategies and prompting for code tasks, adapted from (Wei et al., 2021). For (A), which involves training a Transformer from scratch, please refer to (Ahmad et al., 2020) for its use in source code summarization task. In the case of (E), we utilize a representative RLHF (Ouyang et al., 2022) as an example. Additional reinforcement learning methods, such as DPO (Rafailov et al., 2024), are also applicable at this stage.
@@ 5.4. Instruction Tuning
After pre-training LLMs on large-scale datasets, the next phase typically involves augmenting the model’s ability to process and follow various instructions, known as instruction tuning. Instruction tuning generally refers to the supervised fine-tuning of pre-trained LLMs using datasets comprised of structured examples framed as various natural language instructions (Wei et al., 2021; Ouyang et al., 2022; Iyer et al., 2022; Zhang et al., 2023d). The comparison of instruction tuning with various fine-tuning strategies and prompting for code tasks is depicted in Figure 9. Two exemplars of instruction data sampled from Code Alpaca (Chaudhary, 2023) are demonstrated in Figure 10. It capitalizes on the heterogeneity of instruction types, positioning instruction tuning as a form of multi-task prompted training that significantly enhances the model’s generalization to unseen tasks (Wei et al., 2021; Sanh et al., 2022; Ouyang et al., 2022; Chung et al., 2024).
In the realm of code generation, natural language descriptions serve as the instructions guiding the model to generate corresponding code snippets. Consequently, a line of research on instruction tuning LLMs for code generation has garnered substantial interest across academia and industry. To perform instruction tuning, instruction data are typically compiled from source code with permissive licenses (Husain et al., 2019; Kocetkov et al., 2022; Lozhkov et al., 2024) (refer to Section 5.1.2) or are constructed from synthetic code data (Luo et al., 2023; Wei et al., 2023b; Yuxiang Wei, 2024) (refer to Section 5.2). These datasets are then utilized to fine-tune LLMs through a supervised learning paradigm. However, the substantial computational resources required for full parameter fine-tuning (FFT) LLM pose a notable challenge, particularly in scenarios with constrained resources (Ding et al., 2022a; Lialin et al., 2023). To mitigate this issue, parameter-efficient fine-tuning (PEFT) has emerged as a compelling alternative strategy, gaining increasing attention for its potential to reduce resource consumption (Ding et al., 2022a). In the following subsection, we categorize existing works based on their instruction-tuning strategies to provide a comprehensive and systematic review.
@@ 5.4.1. Full Parameter Fine-tuning
Full parameter fine-tuning (FFT) involves updating all parameters within a pre-trained model, as shown in Figure 11(a). This approach is often preferred when ample computational resources and substantial training data are available, as it typically leads to better performance. (Wang et al., 2021) introduces an encoder-decoder pre-trained language model for code generation, named CodeT5+. They instruction-tune this model on a dataset comprising 20k instruction samples from Code Alpaca (Chaudhary, 2023), resulting in an instruction-following model called InstructCodeT5+, which exhibited improved capabilities in code generation. (Luo et al., 2023) leverages the Evol-Instruct data synthesis technique from WizardLM (Xu et al., 2023) to evolve 20K code Alpaca (Chaudhary, 2023) instruction samples into a 78K code instruction dataset. This enriched dataset is then used to fine-tune the StarCoder base model, resulting in WizardCoder, which showcases notable advancements in code generation. In a similar vein, inspired by the successes of WizardCoder (Luo et al., 2023) and RRHF (Yuan et al., 2023b), Pangu-Coder 2 (Shen et al., 2023) applies the Evol-Instruct method to generate 68k high-quality instruction samples from the initial 20k Code Alpaca (Chaudhary, 2023) instruction samples. Additionally, they introduces a novel reinforcement learning via Rank Responses to align Test & Teacher Feedback (RRTF), which further enhances the performance of Pangu-Coder 2 in code generation. Diverging from synthetic instruction data generation methods, OctoPack (Muennighoff et al., 2023) utilizes real-world data by curating CommitPack from the natural structure of Git commits, which inherently pair code changes with human-written instructions. This dataset, consisting of 4 terabytes of Git commits across 350 programming languages, is employed to fine-tune StarCoder (Li et al., 2023a) and CodeGeeX2 (Zheng et al., 2023c), leading to the instruction-following code models of OctoCoder and OctoGeeX for code generation, respectively. The most recent innovation comes from Magicoder (Wei et al., 2023b), who proposes OSS-INSTRUCT, a novel data synthesis method that leverages open-source code snippets to generate high-quality instruction data for code generation. This approach seeks to reduce the bias often present in synthetic data generated by LLM. In line with OSS-INSTRUCT, the BigCode team introduces StarCoder2-15B-instruct (Yuxiang Wei, 2024), which they claim to be the first entirely self-aligned LLM for code generation, trained with a fully permissive and transparent pipeline. Moreover, (CodeGemma Team et al., 2024) harnesses open-source mathematics datasets, such as MATH (Hendrycks et al., 2021) and GSM8k (Cobbe et al., 2021), along with synthetically generated code following the OSS-INSTRUCT (Wei et al., 2023b) paradigm, to instruction-tune CodeGemma 7B, yielding exceptional results in mathematical reasoning and code generation tasks.
@@ 5.4.2. Parameter-Efficient Fine-tuning
To mitigate the extensive computational and resource demands inherent in fine-tuning LLMs, the concept of parameter-efficient fine-tuning (PEFT) has emerged to focus on updating a minimal subset of parameters, which may either be a selection of the model’s parameters or an array of additional parameters specifically introduced for the tuning process (Ding et al., 2022a; Lialin et al., 2023). The categorization of these methods is depicted in Figure 11(b), (c), and (d). A plethora of innovative PEFT approaches have been developed, among which BitFit (Zaken et al., 2021), Adapter (Houlsby et al., 2019), Prompt tuning (Lester et al., 2021), Prefix-tuning (Li and Liang, 2021), LoRA (Hu et al., 2021), IA3 (Liu et al., 2022a), QLoRA (Dettmers et al., 2024), and AdaLoRA (Zhang et al., 2023a) are particularly noteworthy. A seminal study in this field, LoRA (Hu et al., 2021), proposes a parameter update mechanism for a pre-trained weight matrix — such as those found in the key or value projection matrices of a Transformer block’s multi-head self-attention layer — by factorizing the update into two low-rank matrices. Crucially, all original model parameters remain frozen, with only the pair of low-rank matrices being trainable. After fine-tuning, the product of these low-rank matrices can be seamlessly incorporated into the existing weight matrix through an element-wise addition. This process can be formally described as:
@@
Despite the advancements in PEFT methods, their application in code generation remains limited. For instance, (Jiang and Kim, 2023) pioneered the use of parameter-efficient instruction-tuning on a Llama 2 (Touvron et al., 2023b) model with a single RTX 3090 GPU, leading to the development of a multilingual code generation model called CodeUp. More recently, ASTRAIOS (Zhuo et al., 2024b) conducted a thorough empirical examination of parameter-efficient instruction tuning for code comprehension and generation tasks. This study yielded several perceptive observations and conclusions, contributing valuable insights to the domain.
@@ 5.5. Reinforcement Learning with Feedback
LLMs have exhibited remarkable instruction-following capabilities through instruction tuning. However, they often produce outputs that are unexpected, toxic, biased, or hallucinated outputs that do not align with users’ intentions or preferences (Ouyang et al., 2022; Wang et al., 2023c; Ji et al., 2023b). Consequently, aligning LLMs with human preference has emerged as a pivotal area of research. A notable work is InstructGPT (Ouyang et al., 2022), which further fine-tunes an instruction-tuned model utilizing reinforcement learning with human feedback (RLHF) on a dataset where labelers have ranked model outputs in order of quality, from best to worst. This method has been instrumental in the development of advanced conversational language models, such as ChatGPT (OpenAI, 2022) and Bard (Manyika and Hsiao, 2023). Despite its success, acquiring high-quality human preference ranking data is a resource-intensive process (Lee et al., 2023). To address this, Reinforcement Learning from AI Feedback (RLAIF) (Bai et al., 2022; Lee et al., 2023) has been proposed to leverage powerful off-the-shelf LLMs (e.g., ChatGPT (OpenAI, 2022) and GPT-4 (Achiam et al., 2023)) to simulate human annotators by generating preference data.
@@
Building on RLHF’s success, researchers have explored reinforcement learning with feedback to enhance code generation in LLMs. Unlike RLHF, which relies on human feedback, this approach employs compilers or interpreters to automatically provide feedback on code samples through code execution on unit test cases, catalyzing the advancement of this research domain. CodeRL (Le et al., 2022) introduced an actor-critic reinforcement learning framework for code generation. In this setup, the language model serves as the actor-network, while a token-level functional correctness reward predictor acts as the critic. Generated code is assessed through unit test signals from a compiler, which can indicate compiler errors, runtime errors, unit test failures, or passes. CompCoder (Wang et al., 2022b) enhances code compilability by employing compiler feedback, including language model fine-tuning, compilability reinforcement, and compilability discrimination strategies. Subsequently, PPOCoder (Shojaee et al., 2023) integrates pre-trained code model CodeT5 (Wang et al., 2021) with Proximal Policy Optimization (PPO) (Schulman et al., 2017). This integration not only utilizes execution (i.e., compilers or interpreters) feedback to assess syntactic and functional correctness but also incorporates a reward function that evaluates the syntactic and semantic congruence between abstract syntax tree (AST) sub-trees and data flow graph (DFG) edges in the generated code against the ground truth. Additionally, the framework applies a KL-divergence penalty to maintain fidelity between the actively learned policy and the referenced pre-trained model, enhancing the optimization process. More recently, RLTF (Liu et al., 2023e) has proposed an online reinforcement learning framework that provides fine-grained feedback based on compiler error information and location, along with adaptive feedback that considers the ratio of passed test cases.
@@
Despite these successes, reinforcement learning algorithms face inherent limitations such as inefficiency, instability, extensive resource requirements, and complex hyperparameter tuning, which can impede the performance and scalability of LLMs. To overcome these challenges, recent studies have introduced various variants of RL methods that do not rely on PPO, including DPO (Rafailov et al., 2024), RRHF (Yuan et al., 2023b), and sDPO (Kim et al., 2024). In essence, these methods aim to maximize the likelihood between the logarithm of conditional probabilities of preferred and rejected responses, which may be produced by LLMs with varying capabilities. Inspired by RRHF (Yuan et al., 2023b), PanGu-Coder 2 (Shen et al., 2023) leverages a novel framework, Reinforcement Learning via Rank Responses to align Test & Teacher Feedback (RRTF), significantly enhancing code generation capabilities, as evidenced by pass@1 of 62.20% on the HumanEval benchmark.